### v1

# Symmetries and Anomalies of Hamiltonian Staggered Fermions

Abstract
We review the shift, time reversal and phase symmetries of Hamiltonian staggered fermions and their connection to continuum symmetries concentrating in particular on the case of massless fermions and (3+1) dimensions. We construct operators using the staggered fields that implement these symmetries on finite lattices. We show that it is possible to build shift symmetries that commute with time reversal by combining the elementary shifts with the phase symmetry. In general these symmetries do not form a mutually commuting set and the system suffers from ’t Hooft anomalies. However, in the case where the phase symmetry is broken to $Z_{4}$ by four fermion interactions we find that these lattice anomalies cancel. The naive continuum limit of this anomaly free lattice model has the symmetries and matter representations of the Pati-Salam GUT.
1 Introduction
In this paper, we focus on the symmetries of staggered fermions. Although the symmetries of the Euclidean formulation, in which both time and space are discretized, are well known [1, 2, 3, 4, 5, 6], the Hamiltonian formalism developed in [7] has received less attention in the case when the spatial dimension is greater than one  [^1] Our work has been influenced, however, by recent theoretical work on staggered fermions in (1+1) dimensions given in [8, 9]. In our work we focus on the structure of the shift and time reversal symmetries for Hamiltonian staggered fermions for arbitrary spatial dimension. We are particularly interested in understanding the connection between the anomalies seen in Euclidean formulations of staggered or Kähler-Dirac fermions [10, 11] and the structure of these theories as viewed from a Hamiltonian perspective. In particular, we would like to understand whether we can build chiral lattice gauge theories by gauging certain discrete translation symmetries of staggered fermions along the lines proposed in [12]. These discrete symmetries are called shift symmetries in the literature, and, as we will discuss later, can be thought of as a finite subgroup of the axial-flavor symmetry of the continuum theory. In particular our focus will be on understanding whether such symmetries break in response to gauging other symmetries corresponding to the presence of mixed ’t Hooft anomalies.
Following the procedure given in [13] we construct explicit operators that implement the elementary shifts $S_{k}$, time reversal ${\cal T}$ and a global $U(1)$ symmetry $M$ on a finite lattice. We show how to combine the elementary shifts with a particular element of $M$ to allow the shifts to commute with time reversal. Nevertheless, the group of all symmetries do not form a mutually commuting set which we interpret as signaling the presence of ’t Hooft anomalies [14, 13, 15]. Our work follows on from recent work by on Majorana chains in one spatial dimension [13] and the Schwinger model [9]. Motivation for our work can be found in symmetric mass generation which requires the cancellation of lattice ’t Hooft anomalies [16, 17, 18, 10, 19, 11] and the formulation of certain lattice chiral gauge theories using mirror fermions [20, 21].
We start by reviewing the staggering procedure for the Hamiltonian formalism and then discuss the symmetries of the system focusing on the shift and time reversal symmetries. We then construct finite operators that implement these symmetries on the lattice and their commutator structure. We then focus on the case of three spatial dimensions showing the relationship of shift to continuum flavor symmetries. In this case we show it is possible to cancel potential lattice ’t Hooft anomalies by adding four fermion terms with a particular structure and discuss the continuum limit.
2 Hamiltonian staggered fermions
The continuum Dirac Hamiltonian is given by
$H=\int d^{3}x\,\overline{\Psi}(x,t)\,\left(i\gamma^{i}\partial_{i}+m\right)\Psi(x,t)$ (1)
where $i$ are spatial indices running from $1\ldots d$. The lattice Hamiltonian is obtained by first introducing a cubic spatial lattice and replacing the derivative with a symmetric finite difference
$H=\sum_{x}\overline{\Psi}(x,t)\left[i\gamma^{i}\,(\Delta_{i}(x,y)+m\Psi(x)\right]$ (2)
where $x$ now labels an integer position vector on the lattice with components $\{x_{i}\}$ and the symmetric difference operator is defined by
$\Delta_{i}(x,y)=\frac{1}{2}\left(\delta_{y,x+i}-\delta_{y,x-i}\right)$ (3)
We indicate a shift of the site $x$ by one lattice spacing in the $i^{\rm th}$ direction by $(x+i)$. It is convenient to introduce the hermitian matrices $\alpha_{i}=\gamma_{0}\gamma_{i}$ and $\beta=\gamma_{0}$ and rewrite this as
$H=\sum_{x,y}\Psi^{\dagger}(x,t)\left[\,i\alpha_{i}\Delta_{i}(x,y)+m\beta\delta_{xy}\right]\Psi(y,t)$ (4)
To obtain the staggered Hamiltonian, we first perform a unitary transformation on $\Psi$ to a new basis $\chi$ as follows:
$\Psi(x,t)$ $=\alpha^{x}\chi(x,t)$
$\Psi^{\dagger}(x,t)$ $=\chi^{\dagger}(x,t)(\alpha^{x})^{\dagger}$ (5)
where
$\alpha^{x}=\alpha_{1}^{x_{1}}\alpha_{2}^{x_{2}}\ldots\alpha_{d}^{x_{d}}$ (6)
The Hamiltonian in this basis is then given by
$H=\sum_{x,y}\chi^{\dagger}(x,t)\left[i\eta_{i}(x)\Delta_{i}(x,y)+m\epsilon(x)\delta_{xy}\beta\right]\chi(y,t)$ (7)
where $\eta_{i}(x)=(-1)^{x_{1}+x_{2}+...+x_{i-1}}$ and $\epsilon(x)=\left(-1\right)^{\sum_{i}^{d}x_{i}}$ is the spatial site parity. Unlike the analogous situation in Euclidean space the resultant operator is not proportional to the unit matrix in spinor space. So we cannot stagger the field by discarding all but one component of $\chi$ as one would do in that case. Instead, we can go back to the unitary transformation in eqn. 5 and decompose $\Psi(x)$ into two components $\Psi_{\pm}(x)$ that are eigenstates of $\beta$
$\Psi_{\pm}(x)=P_{\pm}\Psi(x)$ (8)
where $P_{\pm}=\frac{1}{2}\left(1\pm\beta\right)$. Since $\beta$ anticommutes with $\alpha_{i}$ the kinetic operator now couples $\Psi_{+}$ to $\Psi_{-}$. The unitary transformation becomes
$\Psi_{\pm}(x)$ $=P_{\pm}\alpha^{x}\chi(x)=\alpha^{x}\frac{1}{2}(1\pm\beta\epsilon(x))\chi(x)$
$\Psi_{\pm}^{\dagger}(x)$ $=\chi^{\dagger}(x)(\alpha^{x})^{\dagger}P_{\pm}=\chi^{\dagger}(x)\frac{1}{2}(1\pm\beta\epsilon(x))(\alpha^{x})^{\dagger}$ (9)
Thus the fields $\Psi_{+}(x)$ where $\beta=1$ are restricted to even lattice site fields $\chi_{\rm even}(x)=\frac{1}{2}(1+\epsilon(x))\chi(x)$ and the fields $\Psi_{-}(x)$ with $\beta=-1$ to odd site fields $\chi_{\rm odd}(x)$. After truncating $\chi(x)$ to a single component field on each site the final staggered Hamiltonian is then [7, 22, 9, 8]
$H=\sum_{x,y}\left[\chi^{\dagger}(x,t)i\eta_{i}(x)\Delta_{i}(x,y)\chi(y,t)\right]+m\sum_{x}\,\epsilon(x)\chi^{\dagger}(x,t)\chi(x,t)$ (10)
The canonical anticommutators of the staggered fields are given by
$\{\chi^{\dagger}(x,t),\chi(x^{\prime},t)\}=\delta(x,x^{\prime})\quad{\rm with\;all\;other\;anticommutators\;vanishing}$ (11)
while the equation of motion is
$i\frac{\partial\chi(x)}{\partial t}$ $=[H,\chi(x)]$
$=i\eta_{i}(x)\Delta_{i}(x,y)\chi(y)+m\epsilon(x)\chi(x)$ (12)
It is not hard to verify that
$\eta_{i}(x)\eta_{j}(x+i)+\eta_{j}(x)\eta_{i}(x+j)=2\delta_{ij}$ (13)
This result, together with the fact that the site parity operator $\epsilon(x)$ anticommutes with the symmetric difference operator $\Delta_{i}$, ensures that the field $\chi$ satisfies a discrete Klein Gordon equation
$\frac{\partial^{2}\chi(x)}{\partial t^{2}}=\frac{1}{2}\sum_{i}\left[\chi(x+2i)+\chi(x-2i)-2\chi(x)\right]+m^{2}\chi(x)$ (14)
Notice the appearance of a discrete Laplacian operator on a block lattice with twice the lattice spacing.
Let us try to understand how the continuum spinor fields can be reassembled from these staggered fields. For simplicity let us restrict the following discussion to odd $d$. The naive fermion on a d-dimensional spatial lattice gives rise to $2^{d}$ Dirac fermions in the continuum limit because of doubling. After staggering one expects the continuum theory to comprise $2^{d}\times 2^{-(d+1)/2}$ Dirac fermions possessing $2^{d}$ complex components. Each of these components can be identified with the fields $\chi(x)$ living at the corners of a unit cube in the lattice. On a block lattice with twice the lattice spacing we can then build a matrix fermion according to the rule
$\Psi(2x)=\sum_{\{b\}}\chi(x+b)\alpha^{x+b}$ (15)
where $\{b\}$ is a set of $2^{d}$ vectors with components $b_{i}=\{0,1\}$ corresponding to points in the unit cube. The continuum spinors can then be read off from the columns of this matrix as the lattice spacing is sent to zero. However, the most general matrix built from Dirac spinors will possess $2^{\frac{d+1}{2}}\times 2^{\frac{d+1}{2}}=2^{d+1}$ elements. This is twice the number of points in the spatial cube employed in our Hamiltonian construction and will lead to twice as many Dirac fermions in the continuum as that given in eqn. 15. To achieve this doubled set of Dirac fermions one can employ a second staggered fermion $\chi^{\prime}$ and expand it on the additional matrix basis given by $\beta\alpha^{x}$
$\Psi^{\prime}(2x)=\sum_{\{b\}}\chi^{\prime}(x+b)\beta\alpha^{x+b}$ (16)
By adding $\Psi$ and $\Psi^{\prime}$ we can then build a theory containing $2^{\frac{d+1}{2}}$ Dirac fermions. In fact, this is the number of fermions that arises in Euclidean formulations of staggered fermions where time is also discretized. To achieve the number of fermions seen in the minimal Hamiltonian construction comprising just $\Psi$, one needs to project the Euclidean theory by including only Euclidean fermions of a given site parity on the spacetime lattice. This can only be done in the massless case and the resulting formulation is called reduced staggered fermions in the literature - see eg. [2, 18, 21]. Notice that for $d$ odd the matrix $P=i\alpha_{1}\cdots\alpha_{d}=-\gamma_{5}$ commutes with $\Psi$ and anticommutes with $\Psi^{\prime}$. In this case $\Psi$ is an eigenstate of the twisted chiral operator $\gamma_{5}\times\gamma_{5}$ which acts on the left and right of the matrix fermion while $\Psi^{\prime}$ is the other eigenstate.
3 Shift and time reversal symmetries
It is easy to show that the equation of motion eqn. 12 is invariant under time reversal ${\cal T}$:
$t$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-t$
$i$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-i$
$\chi(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\chi^{\dagger}(x)$ (17)
The Hamiltonian is also invariant under a set of translation or shift symmetries $S_{i}$ corresponding to
$\chi(x)\stackrel{{\scriptstyle S_{i}}}{{\to}}i\xi_{i}(x)\chi(x+i)$ (18)
where $\xi_{i}(x)=\left(-1\right)^{\sum_{j=i+1}^{d}x_{j}}$ and we have used the result $\eta_{i}(j)\xi_{j}(i)=1$. The following argument shows that ${\cal T}$ and $S^{k}$ commute.
$\chi(x)$ $\stackrel{{\scriptstyle S_{k}}}{{\to}}i\xi_{k}(x)\chi(x+k)$
$\chi(x)$ $\stackrel{{\scriptstyle{\cal T}\,S_{k}}}{{\to}}-i\xi_{k}(x)\epsilon(x+k)\chi^{*}(x+k)=i\xi_{k}(x)\epsilon(x)\chi^{*}(x+k)$
${\rm similarly}$
$\chi(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\chi^{*}(x)$
$\chi(x)$ $\stackrel{{\scriptstyle S_{k}{\cal T}}}{{\to}}i\xi_{k}(x)\epsilon(x)\chi^{*}(x+k)$ (19)
Furthermore, it should be clear that these simple shifts can be applied consecutively to yield additional symmetries. For example the double shift $S_{ij}$
$\chi(x)\stackrel{{\scriptstyle S_{ij}}}{{\to}}S_{i}S_{j}\,\chi(x)=\xi_{j}(x)\xi_{i}(x+j)\chi(x+i+j)\quad i\neq j$ (20)
It is trivial to see that $S_{i}S_{j}=-S_{j}S_{i}$. Notice that a double shift along the same direction yields a simple translation $T_{i}$ in the i-direction on the block lattice:
$\chi(x)\stackrel{{\scriptstyle S_{i}^{2}}}{{\to}}\xi_{i}(x)\xi_{i}(x+i)\chi(x+2i)=-\chi(x+2i)=-T_{i}\left[\chi(x)\right]$ (21)
Performing a $S_{ij}$ shift followed by a $S_{j}$ shift reveals a non-trivial algebra involving both shifts and translations:
$S_{ij}S_{j}\chi(x)$ $\to S_{ij}i\xi_{j}(x)\chi(x+j)$
$=\xi_{j}(x+j)\xi_{i}(x+2j)\xi_{j}(x)\chi(x+2j+i)=-i\xi_{i}(x)T_{j}\chi(x+i)$
$=-T_{j}S_{i}\chi(x)$ (22)
or more generally $[S_{ij},S_{j}]=-T_{j}S_{i}$. That is, the combination of two shifts generates another shift up to a simple translation on the block lattice. In fact the shift symmetries form a finite group $\Gamma$ whose elements are formed from all possible shifts
$\Gamma=\{\pm I,\pm S_{i},\pm S_{ij},\pm S_{ijk},\cdots,S_{12\ldots d}\}\quad i,j,k=1\ldots d$ (23)
This group closes only on the product of block translations and shifts. The shift symmetries hold even in the presence of gauge interactions provided the gauge link field $U_{i}(x)$ transforms similarly under shifts
$U_{i}(x)\stackrel{{\scriptstyle S_{j}}}{{\to}}U_{i}(x+j)$ (24)
Furthermore, it is straightforward to see that any staggered fermion bilinear formed from products of fields within the hypercube breaks one or more of these shift symmetries. Thus the shifts protect the lattice theory from developing mass terms due to quantum corrections. In this regard they play the same role that continuum axial-flavor symmetries do in preventing fermions from acquiring mass. This is not a coincidence – in fact the shifts correspond to a discrete subgroup of the continuum flavor symmetry as we will discuss later. It is the presence of this discrete subgroup in the lattice theory that guarantees that the full flavor symmetry of the continuum theory is recovered automatically in the continuum limit - see [3, 5, 2, 1, 4] for a discussion of these issues in the context of the Euclidean theory.
Finally, we note that the Hamiltonian is also clearly invariant under a $U(1)$ symmetry $\chi(x)\to e^{i\theta}\chi(x)$ which will play a role in our later discussion.
4 Symmetry operators
To derive explicit operators that implement shifts and time reversal on a finite lattice it is useful to first re-express the staggered fermion $\chi$ in terms of real fields $\lambda^{1}$ and $\lambda^{2}$:
$\chi(x)$ $=\frac{1}{2}\left(\lambda^{1}(x)+i\lambda^{2}(x)\right)$
$\chi^{\dagger}(x)$ $=\frac{1}{2}\left(\lambda^{1}(x)-i\lambda^{2}(x)\right)$
The Hamiltonian is then
$H=\frac{1}{4}\sum_{x,j}\sum_{a=1}^{2}\lambda^{a}(x,t)\,i\eta_{j}\Delta_{j}\,\lambda^{a}(x,t)-im\sum_{x}\epsilon(x)\epsilon^{ab}\lambda^{a}\lambda^{b}$ (26)
and the equal time commutators become
$\{\lambda^{a}(x,t),\lambda^{b}(x^{\prime},t)\}=\delta^{ab}\delta(x,x^{\prime})$ (27)
Clearly, massless staggered fermions give a generalization of the $d=1$ Majorana chain to higher dimensions [23, 13]. The $U(1)$ symmetry discussed earlier yields an $SO(2)$ symmetry acting on the real fields
$\lambda^{a}(x)=M(\theta)^{ab}\lambda^{b}(x)$ (28)
where $M(\theta)=e^{\theta R}$ and $R^{ab}=\epsilon^{ab}=i\sigma_{2}^{ab}$. Similarly, time reversal ${\cal T}$ becomes
$\lambda^{1}(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\lambda^{1}(x)$
$\lambda^{2}(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-\epsilon(x)\lambda^{2}(x)$ (29)
or equivalently $\lambda^{a}(x)\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\sigma_{3}^{ab}\lambda^{b}(x)$. We can write the elementary shift symmetry described in the previous section as $S_{k}=R\hat{S}_{k}$ where $\hat{S}_{k}$ acts only on a single real field $\lambda^{a}(x)$
$\lambda^{a}(x)\stackrel{{\scriptstyle\hat{S}_{k}}}{{\to}}\xi_{k}(x)\lambda^{a}(x+k)$ (30)
When $m=0$, it should be clear that the Hamiltonian is actually invariant under two separate half shifts given by
$\lambda^{1}(x)$ $\stackrel{{\scriptstyle A_{k}}}{{\to}}\xi_{k}(x)\lambda^{1}(x+k)$
$\lambda^{2}(x)$ $\stackrel{{\scriptstyle A_{k}}}{{\to}}\lambda^{2}(x)$ (31)
and
$\lambda^{2}(x)$ $\stackrel{{\scriptstyle B_{k}}}{{\to}}\xi_{k}(x)\lambda^{2}(x+k)$
$\lambda^{1}(x)$ $\stackrel{{\scriptstyle B_{k}}}{{\to}}\lambda^{1}(x)$ (32)
It is not hard to see that the full shifts can be obtained from $A_{k}$ and $B_{k}$
$S_{k}=R\hat{S}_{k}=R\,A_{k}\,B_{k}$ (33)
Unlike the full shifts, the half shifts do not commute with ${\cal T}$ or $M$.
Let us now construct operators that implement $A_{k}$, $B_{k}$ and $\hat{S}_{k}$ on a finite lattice equipped with periodic boundary conditions. As a warm up let us start with a one dimensional lattice with $L$ sites and coordinate $x\equiv x_{1}=0\ldots L-1$. For staggered fermions $L$ must be even and for $d=1$ the phase $\xi_{1}(x)=1$. The $A_{1}$ shift can then be achieved by the action of a shift operator $A_{1}$
$A_{1}=2^{-L/2}\prod^{L-1}_{x=0}\Big{(}1-\lambda^{1}(x)\lambda^{1}(x+1)\Big{)},$ (34)
and $A_{1}^{-1}$ is given by
$A_{1}^{-1}=2^{-L/2}\prod^{x=0}_{L-1}\Big{(}1+\lambda^{1}(x)\lambda^{1}(x+1)\Big{)},$ (35)
To see this one uses the results
$-\lambda^{1}(x+1)$ $=\frac{1}{2}\left[1+\lambda^{1}(x)\lambda^{1}(x+1)\right]\lambda^{1}(x)\left[1-\lambda^{1}(x)\lambda^{1}(x+1)\right]$
$\lambda^{2}(x)$ $=\frac{1}{2}\left[1+\lambda^{1}(x)\lambda^{1}(x+1)\right]\lambda^{2}(x)\left[1-\lambda^{1}(x)\lambda^{1}(x+1)\right]$
11 11 $=\frac{1}{2}\left[1+\lambda^{1}(x)\lambda^{1}(x+1)\right]\left[1-\lambda^{1}(x)\lambda^{1}(x+1)\right]$ (36)
A similar result follows for $B_{1}$ which is given by
$B_{1}=2^{-L/2}\prod^{L-1}_{x=0}\Big{(}1-\lambda^{2}(x)\lambda^{2}(x+1)\Big{)},$ (37)
Combining the $A$ and $B$ shifts one obtains
$\hat{S}_{1}=2^{-L}\prod_{a=1}^{2}\prod^{L-1}_{x=0}\Big{(}1-\lambda^{a}(x)\lambda^{a}(x+1)\Big{)}$ (38)
The time reversal operator can also be implemented in a similar way. When $L$ is a multiple of 4, it can be achieved using the operator ${\cal T}$:
${\cal T}={\cal K}\,\bigg{(}\prod_{x\;{\rm odd}}\lambda^{1}(x)\bigg{)}\bigg{(}\prod_{x\;{\rm even}}\lambda^{2}(x)\bigg{)}.$ (39)
where ${\cal K}$ represents complex conjugation. When $L$ is an odd multiple of 2 this gives the wrong site parity, so we use $G{\cal T}$ instead, where $G$ is the fermion parity operator
$G\lambda^{a}(x)G^{-1}=-\lambda^{a}(x).$ (40)
with $G=\prod_{x}\prod_{a=1}^{2}\lambda^{a}(x)$.
For a two dimensional lattice with coordinates $(x_{1},x_{2})$ with $x_{i}=0\ldots L-1$ the story is similar. An A-shift along $x_{1}$ is given by the action of a shift operator $A_{1}$
$A_{1}=2^{-L^{2}/2}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{1}=0}\Big{(}1-\xi_{1}(x)\lambda^{1}(x_{1},x_{2})\lambda^{1}(x_{1}+1,x_{2})\Big{)}\bigg{)},$ (41)
with $\xi_{1}(x)=\left(-1\right)^{x_{2}}$. Similarly, a shift along $x_{2}$ is generated by the operator
$A_{2}=2^{-L^{2}/2}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{2}=0}\Big{(}1-\xi_{2}(x)\lambda^{1}(x_{1},x_{2})\lambda^{1}(x_{1},x_{2}+1)\Big{)}\bigg{)}.$ (42)
with $\xi_{2}(x)=1$. The B-shifts work in the same way with $\lambda^{1}(x)\to\lambda^{2}(x)$. This allows us to write $\hat{S}_{k}$ take the form
$\hat{S}_{1}$ $=2^{-L^{2}}\prod_{a=1}^{2}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{1}=0}\Big{(}1-\xi_{1}(x)\lambda^{a}(x_{1},x_{2})\lambda^{a}(x_{1}+1,x_{2})\Big{)}\bigg{)}$
$\hat{S}_{2}$ $=2^{-L^{2}}\prod_{a=1}^{2}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{2}=0}\Big{(}1-\xi_{2}(x)\lambda^{a}(x_{1},x_{2})\lambda^{a}(x_{1},x_{2}+1)\Big{)}\bigg{)}$ (43)
When $L$ is a multiple of 4 time reversal can be achieved using
${\cal T}={\cal K}\,\bigg{(}\prod_{x\;{\rm odd}}\lambda^{1}(x_{1},x_{2})\bigg{)}\bigg{(}\prod_{x\;{\rm even}}\lambda^{2}(x_{1},x_{2})\bigg{)},$ (44)
while when $L$ is an odd multiple of 2 we again use $G{\cal T}$ instead. In three dimensions the $\hat{S}$ shifts are
$\hat{S}_{1}$ $=2^{-L^{3}}\prod_{x_{3}=0}^{L-1}\bigg{(}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{1}=0}\Big{(}1-\xi_{1}(x)\lambda^{a}(x_{1},x_{2},x_{3})\lambda^{a}(x_{1}+1,x_{2},x_{3})\Big{)}\bigg{)}\bigg{)}$
$\hat{S}_{2}$ $=2^{-L^{3}}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod_{x_{3}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{2}=0}\Big{(}1-\xi_{2}(x)\lambda^{a}(x_{1},x_{2},x_{3})\lambda^{a}(x_{1},x_{2}+1,x_{3})\Big{)}\bigg{)}\bigg{)}$
$\hat{S}_{3}$ $=2^{-L^{3}}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{3}=0}\Big{(}1-\xi_{3}(x)\lambda^{a}(x_{1},x_{2},x_{3})\lambda^{a}(x_{1},x_{2},x_{3}+1)\Big{)}\bigg{)}\bigg{)}$ (45)
[…]
${\cal T}={\cal K}\,\bigg{(}\prod_{x_{1}+x_{2}+x_{3}={\rm odd}}\lambda^{1}(x_{1},x_{2},x_{3})\bigg{)}\bigg{(}\prod_{x_{1}+x_{2}+x_{3}={\rm even}}\lambda^{2}(x_{1},x_{2},x_{3})\bigg{)}$ (46)
[…]
$R=2^{-L^{d}}\prod_{a=1}^{2}\prod^{L^{d}}_{x=0}\Big{(}1-\epsilon^{ab}\lambda^{a}(x)\lambda^{b}(x)\Big{)}$ (47)
In this way we have constructed explicit operators that implement elementary single shifts $S_{k}$, time reversal ${\cal T}$ and the $R$ symmetry for Hamiltonian staggered fermions. To write down operators that correspond to multiple shifts one simply compounds a series of single shift operators as discussed earlier. It is straightforward to verify the following (anti)commutation relations between the operators $S_{k}$, ${\cal T}$ and $R$:
$[S_{k},{\cal T}]$ $=0$
$[S_{k},R]$ $=0$
$[M,{\cal T}]$ $\neq 0\quad{\rm and\;specifically\;}\{R,{\cal T}\}=0$ (48)
Thus, the symmetries of the system do not constitute a mutually commuting set - the $SO(2)$ rotations do not commute with time reversal. This implies that gauging the global symmetry $M$ will break ${\cal T}$ a ’t Hooft anomaly. Since the odd shifts also depend on $R$ these will also be broken if $M$ is gauged. In the next section we will explore how these ’t anomalies can be cancelled in the interesting case of three dimensions.
5 Three dimensions and anomaly cancellation
As an example, it is instructive to examine the structure of the continuum theory in three dimensions arising from the minimal Hamiltonian theory with one staggered fermion corresponding to the matrix $\Psi$ in eqn. 15. We can adopt a chiral basis for the Dirac gamma matrices
$\gamma_{\mu}=\left(\begin{array}[]{cc}0&\sigma_{\mu}\\ \bar{\sigma}_{\mu}&0\end{array}\right)$ (49)
where $\sigma_{\mu}=(I,\sigma_{i})$ and $\bar{\sigma}_{\mu}=(I,-\sigma_{i})$. In this case
$\Psi=\left(\begin{array}[]{cc}\psi_{R}&0\\ 0&\psi_{L}\end{array}\right)$ (50)
where
$\psi_{R}$ $=\chi(x)I-\chi(x+i)\sigma_{i}-\frac{i}{2}\chi(x+i+j)\epsilon_{ijk}\sigma_{k}+iI\chi(x+1+2+3)$
$\psi_{L}$ $=\chi(x)I+\chi(x+i)\sigma_{i}-\frac{i}{2}\chi(x+i+j)\epsilon_{ijk}\sigma_{k}-iI\chi(x+1+2+3)$ (51)
Clearly, in the naive continuum limit, the staggered field gives rise to two left and two right-handed Weyl fields transforming under two independent $SU(2)$ flavor symmetries. Such a system possesses no invariant bilinear mass terms and hence is chiral  [^2] A Majorana mass for either left or right handed fermions transforming under an $SU(2)$ flavor symmetry vanishes identically.. Of course, at finite lattice spacing there is no exact $SU(2)$ symmetry and one can write down lattice mass terms by combining staggered fields within the unit hypercube. However, such terms break the shift symmetries. As remarked earlier this suggests that the lattice shift symmetries are somewhow related to continuum axial-flavor symmetries. This is indeed the case as we will now show. Flavor rotations of the continuum matrix fermion are given by $e^{i\theta_{A}\alpha_{A}}$ where the hermitian basis $\alpha_{A}$ is given in terms of products of the individual $\alpha_{i}$ matrices:
$\alpha_{A}=\{\alpha_{i},i\alpha_{i}\alpha_{j},i\alpha_{1}\alpha_{2}\alpha_{3}\}\quad{\rm where}\;i=1\ldots 3$ (52)
This yields the flavor group $SU(2)\times SU(2)\times U(1)$ which is the subgroup of $SU(4)$ that commutes with the twisted chiral symmetry $\gamma_{5}\otimes\gamma_{5}$. The $U(1)$ component is generated by $i\alpha_{1}\alpha_{2}\alpha_{3}$ and corresponds to a singlet axial rotation. Under a rotation with generator $\alpha_{j}$ the fermion matrix transforms as
$\Psi\to\Psi e^{i\theta_{j}\alpha_{j}}$ (53)
Staggered fermions, being related to Kähler-Dirac fermions, are only invariant under a twisted rotation group corresponding to the diagonal subgroup of the flavor and rotation symmetries [24, 25, 6, 3, 12]. Upon discretization, the flavor rotation angles are thus further restricted to multiples of $\frac{\pi}{2}$. Then the discrete flavor rotation becomes
$\Psi\to\Psi\,i\alpha_{j}$ (54)
To see that this induces a shift transformation on the staggered fields simply replace the continuum matrix $\Psi$ by its lattice counterpart given in eqn. 15
$\Psi=\sum_{\{b\}}\chi(x+b)\alpha^{x+b}i\alpha_{j}=i\sum_{\{b\}}\chi(x+b)\xi_{j}(x+b)\alpha^{x+b+j}$ (55)
The net effect is clearly to just produce an elementary shift $\chi(x)\to i\xi_{j}(x)\chi(x+j)$. Similar arguments relate the other shift symmetries to other generators of the continuum flavor group. Clearly the algebra of the shifts that was discussed earlier is just inherited from the algebra of these generators. Thus we see that in three dimensions the staggered fermion Hamiltonian is invariant under a set of shift symmetries that correspond to a discrete subgroup of the continuum global symmetry $SU(2)\times SU(2)\times U(1)$. It should be clear that this connection between staggered shifts and continuum symmetries is true in any dimension.
In terms of the real fields the elementary shift is given by
$\lambda^{a}(x)\stackrel{{\scriptstyle S_{j}}}{{\to}}\xi_{j}(x)R^{ab}\lambda^{b}(x+j)$ (56)
Thus shift operations that arise by combining an even number of elementary shifts do not need a factor of $R$ in their definition while shifts using an odd number of elementary shifts require an additional overall factor of $R$. More accurately the shifts $S_{k}$ involve a discrete $Z_{4}$ subgroup of the $M$ transformation corresponding to powers of $R$ with $R^{4}=I$.
$S_{i}$ $=R\hat{S_{i}}$
$S_{ij}$ $=R^{2}\hat{S}_{ij}=-\hat{S}_{ij}$
$S_{123}$ $=R^{3}\hat{S}_{123}=-R\hat{S}_{123}$ (57)
To better understand the anomaly structure of the theory it is helpful to restrict the possible phase rotations to this $Z_{4}$ subgroup. This can be done by the addition of four fermion terms of the form
$G\sum_{x}\chi^{1}(x)\chi^{2}(x)\chi^{3}(x)\chi^{4}(x)+{\rm h.c}$ (58)
which now requires four complex staggered fermions. It also means that the corresponding continuum theory inherits an additional $SU(4)$ flavor symmetry. In this case the fact that ${\cal T}$ anticommutes with $R$ for each staggered field is sufficient to show that the spectrum of the theory will be degenerate as $G\to 0$ (see the appendix). This allows for spontaneous symmetry breaking in the thermodynamic limit. However, it may be possible to avoid this if the four fermion coupling is sufficiently strong that a singlet ground state and a gap can be produced for sufficiently strong coupling $G$. This phenomenon has been termed symmetric mass generation and has been observed in staggered fermion models with four fermion interactions [23, 19, 16, 26, 27, 28, 18, 29]. Such a phase cannot be achieved if there are ’t Hooft anomalies so it is interesting to see how these are cancelled in this case.
To start we need to construct the symmetry operators for four staggered fermions. They take the form of a product of terms for each staggered fermion where the separate factors all commute:
$S_{k}$ $=\prod_{a=1}^{4}R\hat{S}_{k}^{a}$
${\cal T}$ $=\prod_{a=1}^{4}{\cal T}^{a}$
$R$ $=\prod_{a=1}^{4}R^{a}$ (59)
where the superscript labels the staggered fermion $\chi^{a}(x)$ which is acted upon by its own set of operators $R^{a}$, ${\cal T}^{a}$ and $S_{k}^{a}$. It is easy to see that $[R,{\cal T}]=[S_{k},{\cal T}]=[R,S_{k}]=0$. Thus we can gauge the $Z_{4}$ subgroup without breaking the global shift and time reversal symmetries - the lattice theory is free of mixed ’t Hooft anomalies. This agrees with the Euclidean analysis [11]. The absence of lattice anomalies in this theory suggests that the continuum limit is also anomaly free - which is indeed the case as it possesses the global symmetries and matter representations of the Pati-Salam model [30].
6 Conclusions
In this paper we have examined the shift, time reversal and phase symmetries of Hamiltonian staggered fermions on finite $d$-dimensional spatial lattices. In particular, we have shown how the shift symmetries correspond to a discrete subgroup of the product of the continuum flavor symmetry and translations and play a crucial role in protecting the theory from developing lattice mass terms as a result of quantum corrections. We have constructed explicit operators to generate these symmetries along the lines of [13] with the massless staggered fermions describing an analog of Majorana chains in more than one dimension. In general the lattice theory possesses ’t Hooft anomalies corresponding to non-vanishing commutators between subsets of these symmetries but we have shown that these can be canceled in the presence of four fermion terms of a particular type.
While writing this paper we became aware of another recent work which also elucidates the symmetry structure of Hamiltonian staggered fermions with the goal of classifying the possibilities for symmetric mass generation [31]. Our results are consistent with their conclusions where the two papers overlap.
Acknowledgements
This work was supported by the US Department of Energy (DOE), Office of Science, Office of High Energy Physics, under Award Number DE-SC0009998.
Appendix A Energy level degeneracy
It is straightforward to see that the two hermitian symmetry operators that anticommute implies degenerate energy levels. Consider two operators that commute with the Hamiltonian but anticommute with each other
$[H,A]=0\quad[H,B]=0\quad{\rm with}\quad\{A,B\}=0$ (60)
The eigenvectors of $H$ can be labeled with say the eigenvalues $\lambda_{A}$ of $A$
$H\ket{\lambda_{A}}=E\ket{\lambda_{A}}$ (61)
Acting on left with $B$ we find that
$H\,B\ket{\lambda_{A}}=E\,B\ket{\lambda_{A}}$ (62)
Thus the state $B\ket{\lambda_{A}}$ also has energy $E$. To show that it is not proportional to |$\ket{\lambda_{A}}$ consider
$\bra{\lambda_{A}}(AB+BA)\ket{\lambda_{A}}=0$ (63)
But this implies
$\lambda_{A}\braket{\lambda_{A}}{B\lambda_{A}}=0$ (64)
Hence |$\ket{\lambda_{A}}$ and $B\ket{\lambda_{A}}$ are orthogonal for $\lambda_{A}\neq 0$.