### v3

# Symmetries and Anomalies of Hamiltonian Staggered Fermions

Abstract
We review the shift (translation) and time reversal symmetries of Hamiltonian staggered fermions and their connection to continuum symmetries concentrating in particular on the case of massless fermions and (3+1) dimensions. We construct operators using the staggered fields that implement these symmetries on finite lattices. We show that shifts composed of an odd multiple of the elementary shift anti-commute with time reversal and are related to continuum axial transformations. We argue that the presence of these non-trivial commutators implies the existence of exact lattice ’t Hooft anomalies.
preprint:
I Introduction
In this paper, we focus on the symmetries of staggered fermions. Although the symmetries of the Euclidean formulation, in which both time and space are discretized, are well known [1, 2, 3, 4, 5, 6], the Hamiltonian formalism developed in [7] has received less attention in the case when the spatial dimension is greater than one  [^1] Our work has been influenced, however, by recent theoretical work on staggered fermions in (1+1) dimensions given in [8, 9]. In our work we focus on the structure of the shift and time reversal symmetries for Hamiltonian staggered fermions for arbitrary spatial dimension. We are particularly interested in understanding the connection between the anomalies seen in Euclidean formulations of staggered or Kähler-Dirac fermions [10, 11] and the structure of these theories as viewed from a Hamiltonian perspective. In particular, we would like to understand whether we can build chiral lattice gauge theories by gauging certain discrete translation symmetries of staggered fermions along the lines proposed in [12]. These discrete symmetries are called shift symmetries in the literature, and, as we will discuss later, can be thought of as a finite subgroup of the axial-flavor symmetry of the continuum theory. In particular our focus will be on understanding whether such symmetries break in response to gauging other symmetries signaling the presence of mixed ’t Hooft anomalies - a phenomenon that has has been observed in other lattice systems [13, 14, 15].
Following the procedure given in [14] we construct explicit operators that implement the elementary shifts $S_{k}$, time reversal ${\cal T}$ and a global $U(1)$ symmetry $M$ on a finite lattice. Our work can be seen as extension of recent work on Majorana chains in one spatial dimension [14] and the Schwinger model [8, 9]. Motivation for our work can also be found in the phenomenon of symmetric mass generation which requires the cancellation of lattice ’t Hooft anomalies [16, 17, 18, 10, 19, 11] and the formulation of certain lattice chiral gauge theories using mirror fermions [20, 21].
We start by reviewing the staggering procedure for the Hamiltonian formalism and then discuss the symmetries of the system focusing on the shift and time reversal symmetries. We then construct finite operators that implement these symmetries on the lattice and examine their commutator structure. We examine in particular the case of three spatial dimensions showing the relationship of shift invariances to continuum axial-flavor symmetries.
II Hamiltonian staggered fermions
The continuum Dirac Hamiltonian is given by
$H=\int d^{3}x\,\overline{\Psi}(x,t)\,\left(i\gamma^{i}\partial_{i}+m\right)\Psi(x,t)$ (1)
where $i$ are spatial indices running from $1\ldots d$. The lattice Hamiltonian is obtained by first introducing a cubic spatial lattice and replacing the derivative with a symmetric finite difference
$H=\sum_{x}\overline{\Psi}(x,t)\left[i\gamma^{i}\,(\Delta_{i}(x,y)+m\right]\Psi(y,t)$ (2)
where $x$ now labels an integer position vector on the lattice with components $\{x_{i}\}$ and the symmetric difference operator is defined by
$\Delta_{i}(x,y)=\frac{1}{2}\left(\delta_{y,x+i}-\delta_{y,x-i}\right)$ (3)
We indicate a shift of the site $x$ by one lattice spacing in the $i^{\rm th}$ direction by $(x+i)$. It is convenient to introduce the hermitian matrices $\alpha_{i}=\gamma_{0}\gamma_{i}$ and $\beta=\gamma_{0}$ and rewrite this as
$H=\sum_{x,y}\Psi^{\dagger}(x,t)\left[\,i\alpha_{i}\Delta_{i}(x,y)+m\beta\delta_{xy}\right]\Psi(y,t)$ (4)
To obtain the staggered Hamiltonian, we first perform a unitary transformation on $\Psi$ to a new basis $\chi$ as follows:
$\Psi(x,t)$ $=\alpha^{x}\chi(x,t)$
$\Psi^{\dagger}(x,t)$ $=\chi^{\dagger}(x,t)(\alpha^{x})^{\dagger}$ (5)
where
$\alpha^{x}=\alpha_{1}^{x_{1}}\alpha_{2}^{x_{2}}\ldots\alpha_{d}^{x_{d}}$ (6)
The Hamiltonian in this basis is then given by
$H=\sum_{x,y}\chi^{\dagger}(x,t)\left[i\eta_{i}(x)\Delta_{i}(x,y)+m\epsilon(x)\delta_{xy}\beta\right]\chi(y,t)$ (7)
where $\eta_{i}(x)=(-1)^{x_{1}+x_{2}+...+x_{i-1}}$ and $\epsilon(x)=\left(-1\right)^{\sum_{i}^{d}x_{i}}$ is the spatial site parity. Unlike the analogous situation in Euclidean space the resultant operator is not proportional to the unit matrix in spinor space. So we cannot stagger the field by merely discarding all but one component of $\chi$ as one would do in that case. Instead, we can go back to the unitary transformation in eqn. II and decompose $\Psi(x)$ into two components $\Psi_{\pm}(x)$ that are eigenstates of $\beta$
$\Psi(x)=\Psi_{+}(x)+\Psi_{-}(x)$ (8)
where $P_{\pm}=\frac{1}{2}\left(1\pm\beta\right)$. The unitary transformation becomes
$\Psi_{\pm}(x)=P_{\pm}\alpha^{x}\chi(x)=\alpha^{x}\frac{1}{2}(1\pm\beta\epsilon(x))\chi(x)$ (9)
Since $\beta$ anticommutes with $\alpha_{i}$ the kinetic operator now couples $\Psi^{\dagger}_{+}$ to $\Psi_{-}$ and vice versa. The mass term is diagonal in this basis
$\Psi^{\dagger}\beta\Psi=\Psi_{+}^{\dagger}\Psi_{+}-\Psi_{-}^{\dagger}\Psi_{-}$ (10)
The fields $\Psi_{+}(x)$ where $\beta=1$ are hence restricted to even lattice sites $\chi_{\rm even}(x)=\frac{1}{2}(1+\epsilon(x))\chi(x)$ and the fields $\Psi_{-}(x)$ to odd sites $\chi_{\rm odd}(x)=\frac{1}{2}(1-\epsilon(x))\chi(x)$. The mass term becomes
$\epsilon(x)\Psi^{\dagger}(x)\Psi(x)$ (11)
and we can now truncate $\chi(x)$ to a single component field on each site with the final staggered Hamiltonian  [^2] In this expression $\chi^{\dagger}$ is a single component field and is hence equal to $\chi^{*}$. being given by [7, 22, 9, 8]
$H=$ $\sum_{x,y}\chi^{\dagger}(x,t)i\eta_{i}(x)\Delta_{i}(x,y)\chi(y,t)$
$+\;m\sum_{x}\,\epsilon(x)\chi^{\dagger}(x,t)\chi(x,t)$ (12)
The canonical anticommutators of the staggered fields are given by
$\{\chi^{\dagger}(x,t),\chi(x^{\prime},t)\}=\delta(x,x^{\prime})$ (13)
$\quad{\rm with\;all\;other\;anticommutators\;vanishing}$
while the equation of motion is
$i\frac{\partial\chi(x)}{\partial t}$ $=[H,\chi(x)]$
$=i\eta_{i}(x)\Delta_{i}(x,y)\chi(y)+m\epsilon(x)\chi(x)$ (14)
It is not hard to verify that
$\eta_{i}(x)\eta_{j}(x+i)+\eta_{j}(x)\eta_{i}(x+j)=2\delta_{ij}$ (15)
This result, together with the fact that the site parity operator $\epsilon(x)$ anticommutes with the symmetric difference operator $\Delta_{i}$, ensures that the field $\chi$ satisfies a discrete Klein Gordon equation
$\frac{\partial^{2}\chi(x)}{\partial t^{2}}=\frac{1}{2}\sum_{i}\left[\chi(x+2i)+\chi(x-2i)-2\chi(x)\right]+m^{2}\chi(x)$ (16)
Notice the appearance of a discrete Laplacian operator on a block lattice with twice the lattice spacing. This equation implies that there are $2^{d}$ degenerate solutions for every site on the block lattice. These solutions can be constructed by performing a translation of the block lattice solution to every site in the unit hypercube of the original lattice.
Let us try to understand how this works in more detail. For simplicity let us restrict the following discussion to odd $d$. The naive fermion on a d-dimensional spatial lattice gives rise to $2^{d}$ Dirac fermions in the continuum limit because of doubling. After staggering one expects the continuum theory to correspond to $2^{\frac{d-1}{2}}$ Dirac fermions with a total of $2^{d}$ complex components. These can be identified with the staggered fields $\chi(x)$ living at the corners of the unit hypercube in the original lattice or equivalently the unit cell in the block lattice. On the block lattice we can then build a matrix fermion according to the rule
$\Psi(2x)=\sum_{\{b\}}\chi(x+b)\alpha^{x+b}$ (17)
where $\{b\}$ is a set of $2^{d}$ vectors with components $b_{i}=\{0,1\}$ corresponding to points in the unit cell of the block lattice. The continuum spinors can then be read off from the columns of this matrix as the lattice spacing is sent to zero.
Since a single Dirac fermion has $2^{\frac{d+1}{2}}$ components the most general square matrix built from such spinors will possess $2^{\frac{d+1}{2}}\times 2^{\frac{d+1}{2}}=2\times 2^{d}$ elements. This is twice the number of points in the spatial cube employed in our Hamiltonian construction. To achieve this doubled set of Dirac fermions one can employ a second staggered fermion $\chi^{\prime}$ and expand it on the additional matrix basis given by $\beta\alpha^{x}$
$\Psi^{\prime}(2x)=\sum_{\{b\}}\chi^{\prime}(x+b)\beta\alpha^{x+b}$ (18)
By adding $\Psi$ and $\Psi^{\prime}$ we can then build a theory containing $2^{\frac{d+1}{2}}$ Dirac fermions. In fact, this is the number of fermions that arises in Euclidean formulations of staggered fermions where time is also discretized. To achieve the number of fermions seen in the minimal Hamiltonian construction comprising just $\Psi$, one needs to project the Euclidean theory by including only Euclidean fermions of a given site parity on the spacetime lattice. This can only be done in the massless case and the resulting formulation is called reduced staggered fermions in the literature - see eg. [2, 18, 21]. Notice that for $d$ odd the matrix $P=i\alpha_{1}\cdots\alpha_{d}=\gamma_{5}$ commutes with $\Psi$ and anticommutes with $\Psi^{\prime}$. Thus $\Psi$ is an eigenstate of the twisted chiral operator $\gamma_{5}\otimes\gamma_{5}$ which acts on the left and right of the matrix fermion with eigenvalue $+1$ while $\Psi^{\prime}$ is the other eigenstate with eigenvalue $-1$.
Finally, we note that the Hamiltonian is also clearly invariant under a $U(1)$ symmetry $M$ in which $\chi(x)\to e^{i\theta}\chi(x)$ which will play an important role in our later discussion. We now turn to a discussion of important additional symmetries of the staggered Hamiltonian.
III Continuum flavor and lattice shift symmetries
In this section we will focus on staggered fermions in $3+1$ dimensions but the arguments can be easily generalized to any dimension. We start from a chiral basis for the Dirac gamma matrices given by
$\gamma_{\mu}=\left(\begin{array}[]{cc}0&\sigma_{\mu}\\ \bar{\sigma}_{\mu}&0\end{array}\right)$ (19)
where $\sigma_{\mu}=(I,\sigma_{i})$ and $\bar{\sigma}_{\mu}=(I,-\sigma_{i})$. In this case
$\alpha_{i}=\left(\begin{array}[]{cc}-\sigma_{i}&0\\ 0&\sigma_{i}\end{array}\right)$ (22)
and the matrix $\Psi$ in eqn. 17 then takes the form
$\Psi=\left(\begin{array}[]{cc}\psi_{R}&0\\ 0&\psi_{L}\end{array}\right)$ (24)
where
$\psi_{R}$ $=\chi(x)I-\chi(x+i)\sigma_{i}-\frac{i}{2}\chi(x+i+j)\epsilon_{ijk}\sigma_{k}$
$+iI\chi(x+1+2+3)$
$\psi_{L}$ $=\chi(x)I+\chi(x+i)\sigma_{i}-\frac{i}{2}\chi(x+i+j)\epsilon_{ijk}\sigma_{k}$
$-iI\chi(x+1+2+3)$ (25)
In the naive continuum limit the massless staggered field thus gives rise to doublets of left and right-handed Weyl fields transforming under two independent $SU(2)$ symmetries. These symmetries protect the effective action of the theory from developing bilinear mass terms  [^3] A Majorana mass for either left or right handed doublet vanishes identically . In addition the theory is invariant under both the vector $U(1)$ symmetry discussed earlier and a singlet axial symmetry $U(1)_{A}$ in which the $\psi_{R}$ and $\psi_{L}$ carry opposite charges.
Axial-flavor rotations act by right multiplication of the continuum matrix fermion given in eqn. 24 by the matrix $e^{i\theta_{A}\alpha_{A}}$ where the hermitian basis $\alpha_{A}$ is given in terms of products of the individual $\alpha_{i}$ matrices:
$\alpha_{A}=\{\alpha_{i},i\alpha_{i}\alpha_{j},i\alpha_{1}\alpha_{2}\alpha_{3}\}\quad{\rm where}\;i=1\ldots 3$ (26)
It can be seen that this yields the flavor group $SU(2)\times SU(2)\times U_{A}(1)$ as described above.
Staggered fermions, being discretizations of Kähler-Dirac fermions [11], are invariant under a twisted rotation group corresponding to the diagonal subgroup of the flavor and rotation symmetries [23, 24, 6, 3, 12]. Upon discretization the rotational symmetries are restricted to the cubic group and hence the flavor rotation angles are similarly restricted to odd multiples of $\frac{\pi}{2}$. Then the remaining elementary discrete flavor rotation becomes
$\Psi\to\Psi e^{i\frac{\pi}{2}\alpha_{j}}=\Psi\,i\alpha_{j}$ (27)
We will now show that this induces a unit translation or shift on the staggered fields. Acting on the lattice fermion matrix given in eqn. 17 one finds
$\Psi(2x)$ =$=\sum_{\{b\}}\chi(x+b)\alpha^{x+b}i\alpha_{j}$
$=i\sum_{\{b\}}\xi_{j}(x+b)\chi(x+b)\alpha^{x+b+j}$ (28)
where one must anticommute the $\alpha_{j}$ matrix from the right which produces the phase factor $\xi_{j}(x+b)$. The net effect is clearly to just produce an elementary shift
$\chi(x)$ $\stackrel{{\scriptstyle S_{j}}}{{\to}}i\xi_{j}(x)\chi(x+j)$
$\chi^{\dagger}(x)$ $\stackrel{{\scriptstyle S_{j}}}{{\to}}-i\xi_{j}(x)\chi^{\dagger}(x+j)$ (29)
where $\xi_{j}(x)=\left(-1\right)^{\sum_{k=j+1}^{d}x_{k}}$. It is straightforward to examine the invariance of the Hamiltonian under such a shift $S_{k}$
$H\to$ $\sum_{x,i}i\eta_{i}(x)\chi^{\dagger}(x+k)\xi_{k}(x)\xi_{k}(x+i)\frac{1}{2}\left[\chi(x+i+k)-\right.$
$\left.\quad\chi(x-i+k)\right]+\sum_{x}m\epsilon(x)\xi_{k}(x)^{2}\chi^{\dagger}(x+k)\chi(x+k)$
$\to$ $\sum_{x,i}i\eta_{i}(x-k)\chi^{\dagger}(x)\xi_{k}(x-k)\xi_{k}(x+i-k)\frac{1}{2}\left[\chi(x+i)-\right.$
$\left.\quad\chi(x-i)\right]+\sum_{x}m\epsilon(x-k)\chi^{\dagger}(x)\chi(x)$
$\to$ $\sum_{x}i\eta(x)\chi^{\dagger}(x)\frac{1}{2}\left[\chi(x+i)-\chi(x-i)\right]$
$-\sum_{x}m\epsilon(x)\chi^{\dagger}(x)\chi(x)$ (30)
where we have used the result $\eta_{i}(k)\xi_{k}(i)=1$. Notice that the Hamiltonian is only invariant under $S_{k}$ if $m=0$. It should be clear that elementary shifts can be applied consecutively to yield additional symmetries. For example the double shift $S_{ij}$:
$\chi(x)\stackrel{{\scriptstyle S_{ij}}}{{\to}}S_{i}S_{j}\,\chi(x)=-\xi_{j}(x)\xi_{i}(x+j)\chi(x+i+j)\quad i\neq j$ (31)
It is trivial to see that $S_{i}S_{j}=-S_{j}S_{i}$. Thus, just as $S_{k}$ is associated with the generator $\alpha_{k}$ the properties of the double shift $S_{ij}$ mimic those of the generator $i\alpha_{i}\alpha_{j}$. It is important to notice that in the case of the double shift the Hamiltonian is invariant even for non-zero mass. This is consistent with the vector character of the transformations induced by $\alpha_{i}\alpha_{j}$ and we can conclude that any even number of shifts will correspond to vector symmetries of the continuum theory. Conversely the Hamiltonian is only invariant under an odd shift if the mass is zero. An explicit example of this is the the triple shift $S_{1}S_{2}S_{3}$ which corresponds to the generator of the axial $U_{A}(1)$ symmetry $\gamma_{5}=i\alpha_{1}\alpha_{2}\alpha_{3}$.
It is important to recognize that the connection between lattice shifts and continuum flavor is only one to one up to ordinary translations. For example, a double shift along the same direction yields a simple translation $T$ on the block lattice:
$\chi(x)\stackrel{{\scriptstyle S_{i}^{2}}}{{\to}}-\xi_{i}(x)\xi_{i}(x+i)\chi(x+2i)=\chi(x+2i)=-T_{i}\left[\chi(x)\right]$ (32)
Similarly, performing a $S_{ij}$ shift followed by a $S_{j}$ shift yields
$S_{ij}S_{j}\chi(x)$ $\to S_{ij}i\xi_{j}(x)\chi(x+j)$
$=-i\xi_{j}(x+j)\xi_{i}(x+2j)\xi_{j}(x)\chi(x+2j+i)$
$=-i\xi_{i}(x)T_{j}\chi(x+i)$
$=-T_{j}S_{i}\chi(x)$ (33)
or more generally $[S_{ij},S_{j}]=-2T_{j}S_{i}$. That is, the combination of two shifts generates another shift up to a block translation. In a similar fashion the double shifts satisfy the relation
$[S_{ij},S_{jk}]=2S_{ik}T_{j}$ (34)
It should now be clear that the staggered fermion shift symmetries form a discrete subgroup of the continuum symmetries and guarantee that the latter is restored automatically in the continuum limit. Furthermore, the odd shifts mimic the properties of the continuum axial symmetries in protecting the lattice theory from mass terms constructed from products of the staggered field within the unit hypercube - see [3, 5, 2, 1, 4] for a discussion of these issues in the context of the Euclidean theory. The shift symmetries hold even in the presence of gauge interactions provided the gauge link field $U_{i}(x)$ transforms similarly under shifts
$U_{i}(x)\stackrel{{\scriptstyle S_{j}}}{{\to}}U_{i}(x+j)$ (35)
IV Lattice axial charge
We have already remarked that a discrete $U_{A}(1)$ transformation corresponds to right multiplication of the matrix fermion by $\gamma_{5}$ and maps to the 3-shift $S_{123}$ operator acting on the staggered field. Its axial character can be made obvious by observing that
$\Psi\gamma_{5}=(\gamma_{5}\Psi\gamma_{5})\gamma_{5}=\gamma_{5}\Psi$ (36)
where we have used the fact that the staggered field has eigenvalue unity under the twisted chiral operator $\gamma_{5}\otimes\gamma_{5}$. A straightforward expression for the singlet axial charge can then be written as
$Q_{A}$ $=\frac{1}{2}\sum_{x}\Xi_{123}(x)\chi^{*}(x)\chi(x+1+2+3)+{\rm h.c}$
$=\frac{1}{2}\sum_{x}\Xi_{123}(x)\chi^{*}(x)\left[\chi(x+1+2+3)\right.$
$\left.\quad+\chi(x-1-2-3)\right]$ (37)
where $\Xi_{123}(x)=\xi_{1}(x)\xi_{2}(x+1)\xi_{3}(x+1+2+3)$. This generalizes the result for the non-quantized axial charge $\tilde{Q}_{A}$ given in [9] for one (spatial) dimension to three dimensions. It has a vanishing commutator with the vector charge $Q_{V}=\frac{1}{2}\sum_{x}(\chi^{*}(x)\chi(x)-1/2)$. Following [9] we can also introduce a quantized axial charge using
$Q_{A}$ $=\frac{1}{2}\sum_{x}\Xi_{123}(x)\left[\chi(x)+\chi^{*}(x)\right]\left[\chi(x+1+2+3)\right.$
$\left.\quad-\chi^{*}(x+1+2+3)\right]$ (38)
In this case one finds
$[Q_{V},Q_{A}]$ $=-\frac{1}{2}\sum_{x}\Xi_{123}(x)\left[\chi(x)\chi(x+1+2+3)\right.$
$\left.\quad+\chi^{*}(x)\chi^{*}(x+1+2+3)\right]$ (39)
This is clearly non-zero on a finite lattice and suggests that the theory suffers from a mixed ’t Hooft anomaly that prevents gauging the vector symmetry. One might imagine constructing other charges corresponding to the full set of shift symmetries and examine their commutator structure too in the search for lattice anomalies. However, this is problematic since the shifts correspond to finite rather than infinitessimal transformations in the continuum theory.
V Time reversal on and off the lattice
In (3+1) dimensions the discrete symmetries of charge conjugation $C$ and time reversal $T$ act on the continuum fermion field in the following manner:
$\Psi\stackrel{{\scriptstyle C}}{{\to}}\gamma_{2}\Psi^{*}=\beta\alpha_{2}\Psi^{*}\$ (40)
and
$\Psi$ $\stackrel{{\scriptstyle T}}{{\to}}\gamma_{1}\gamma_{3}\Psi=\alpha_{1}\alpha_{3}\Psi$
$t$ $\stackrel{{\scriptstyle T}}{{\to}}-t$
$i$ $\stackrel{{\scriptstyle T}}{{\to}}-i$ (41)
In particular the combination ${\cal T}=CT$ acts as follows
$\Psi$ $\to\stackrel{{\scriptstyle{\cal T}}}{{\to}}A\Psi^{*}$
$t$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-t$
$i$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-i$ (42)
where $A=\beta\alpha_{2}\alpha_{1}\alpha_{3}$. A standard calculation shows that if $\Psi(x,t)$ is a solution of the EOM then so is $A\Psi^{*}(-t,x)$. Performing the $CT$ transformation on the Hamiltonian after the unitary transformation given in eqn. II shows that the symmetry operation ${\cal T}$ acts on staggered fields as follows
$\chi(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\chi^{*}(x)$ (43)
$t$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-t$
$i$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-i$
For brevity we will refer to this product of charge conjugation and time reversal symmetry as simply time reversal ${\cal T}$ in the rest of the paper  [^4] Notice that ${\cal T}^{2}=1$.. At this point it is important to notice that the elementary shift symmetry does not commute with time reversal:
$\chi(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\chi^{*}(x)$
$\chi(x)$ $\stackrel{{\scriptstyle S_{k}{\cal T}}}{{\to}}-i\epsilon(x)\xi_{k}(x)\chi^{*}(x+k)$
$\chi(x)$ $\stackrel{{\scriptstyle S_{k}}}{{\to}}i\xi_{k}(x)\chi(x+k)$
$\chi(x)$ $\stackrel{{\scriptstyle{\cal T}S_{k}}}{{\to}}-i\xi_{k}(x)\epsilon(x+k)\chi^{*}(x+k)=i\epsilon(x)\xi_{k}(x)\chi^{*}(x+k)$ (44)
Clearly, the two symmetries anticommute. Indeed, this statement is true for any shift composed of an odd number of elementary shifts.
Let us summarise our conclusions so far. The massless staggered theory is invariant under a $U(1)$ phase symmetry, time reversal invariance and a set of shift symmetries that are related to axial-flavor symmetries of the continuum Dirac fields. However, we find that odd multiples of the elementary shift anticommute with time reversal.
VI Real fields and symmetry operators
To derive explicit operators that implement shifts and time reversal on a finite lattice it is useful to first re-express the staggered fermion $\chi$ in terms of real fields $\lambda^{1}$ and $\lambda^{2}$:
$\chi(x)$ $=\frac{1}{2}\left(\lambda^{1}(x)+i\lambda^{2}(x)\right)$
$\chi^{\dagger}(x)$ $=\frac{1}{2}\left(\lambda^{1}(x)-i\lambda^{2}(x)\right)$
The massless Hamiltonian is then
$H=\frac{1}{4}\sum_{x,j}\sum_{a=1}^{2}\lambda^{a}(x,t)\,i\eta_{j}\Delta_{j}\,\lambda^{a}(x,t)$ (46)
and the equal time commutators become
$\{\lambda^{a}(x,t),\lambda^{b}(x^{\prime},t)\}=\delta^{ab}\delta(x,x^{\prime})$ (47)
The $U(1)$ symmetry discussed earlier yields an $SO(2)$ symmetry acting on the real fields
$\lambda^{a}(x)=M(\theta)^{ab}\lambda^{b}(x)$ (48)
where $M(\theta)=e^{\theta R}$ and $R^{ab}=\epsilon^{ab}$. Similarly, time reversal ${\cal T}$ acts as
$\lambda^{1}(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}\epsilon(x)\lambda^{1}(x)$
$\lambda^{2}(x)$ $\stackrel{{\scriptstyle{\cal T}}}{{\to}}-\epsilon(x)\lambda^{2}(x)$ (49)
We can write the elementary shift symmetry described in the previous section as $S_{k}=R\hat{S}_{k}$ where $\hat{S}_{k}$ is given by
$\lambda^{a}(x)\stackrel{{\scriptstyle\hat{S}_{k}}}{{\to}}\xi_{k}(x)\lambda^{a}(x+k)$ (50)
In fact, it should be clear that the massless Hamiltonian is actually invariant under two separate half shifts given by
$\lambda^{1}(x)$ $\stackrel{{\scriptstyle A_{k}}}{{\to}}\xi_{k}(x)\lambda^{1}(x+k)$
$\lambda^{2}(x)$ $\stackrel{{\scriptstyle A_{k}}}{{\to}}\lambda^{2}(x)$ (51)
and
$\lambda^{2}(x)$ $\stackrel{{\scriptstyle B_{k}}}{{\to}}\xi_{k}(x)\lambda^{2}(x+k)$
$\lambda^{1}(x)$ $\stackrel{{\scriptstyle B_{k}}}{{\to}}\lambda^{1}(x)$ (52)
with
$\hat{S}_{k}=A_{k}\,B_{k}$ (53)
The $B$ half shift is precisely the same symmetry considered in [9] for a (1+1) dimensional staggered fermion model.
Let us now construct operators that implement $A_{k}$, $B_{k}$ and $R$ on a finite lattice equipped with periodic boundary conditions. As a warm up let us start with a one dimensional lattice with $L$ sites and coordinate $x\equiv x_{1}=0\ldots L-1$. For staggered fermions $L$ must be even and for $d=1$ the phase $\xi_{1}(x)=1$. The $A_{1}$ shift can then be achieved by the action of a shift operator $A_{1}$
$A_{1}=2^{-L/2}\prod^{L-1}_{x=0}\Big{(}1-\lambda^{1}(x)\lambda^{1}(x+1)\Big{)},$ (54)
and $A_{1}^{-1}$ is given by
$A_{1}^{-1}=2^{-L/2}\prod^{x=0}_{L-1}\Big{(}1+\lambda^{1}(x)\lambda^{1}(x+1)\Big{)},$ (55)
To see this one uses the results
$-\lambda^{1}(x+1)$ $=\frac{1}{2}\left[1+\lambda^{1}(x)\lambda^{1}(x+1)\right]\lambda^{1}(x)\left[1-\right.$
$\left.\quad\lambda^{1}(x)\lambda^{1}(x+1)\right]$
$\lambda^{2}(x)$ $=\frac{1}{2}\left[1+\lambda^{1}(x)\lambda^{1}(x+1)\right]\lambda^{2}(x)\left[1-\right.$
$\left.\quad\lambda^{1}(x)\lambda^{1}(x+1)\right]$
11 11 $=\frac{1}{2}\left[1+\lambda^{1}(x)\lambda^{1}(x+1)\right]\left[1-\right.$
$\left.\quad\lambda^{1}(x)\lambda^{1}(x+1)\right]$ (56)
A similar result follows for $B_{1}$ which is given by
$B_{1}=2^{-L/2}\prod^{L-1}_{x=0}\Big{(}1-\lambda^{2}(x)\lambda^{2}(x+1)\Big{)},$ (57)
Combining the $A$ and $B$ shifts one obtains
$\hat{S}_{1}=2^{-L}\prod_{a=1}^{2}\prod^{L-1}_{x=0}\Big{(}1-\lambda^{a}(x)\lambda^{a}(x+1)\Big{)}$ (58)
The time reversal operator can also be implemented in a similar way. When $L$ is a multiple of 4, it can be achieved using the operator ${\cal T}$:
${\cal T}={\cal K}\,\bigg{(}\prod_{x\;{\rm even}}\lambda^{1}(x)\bigg{)}\bigg{(}\prod_{x\;{\rm odd}}\lambda^{2}(x)\bigg{)}.$ (59)
where ${\cal K}$ represents complex conjugation. When $L$ is an odd multiple of 2 this gives the wrong site parity, so we use $G{\cal T}$ instead, where $G$ is the fermion parity operator
$G\lambda^{a}(x)G^{-1}=-\lambda^{a}(x).$ (60)
with $G=\prod_{x}\prod_{a=1}^{2}\lambda^{a}(x)$.
For a two dimensional lattice with coordinates $(x_{1},x_{2})$ with $x_{i}=0\ldots L-1$ the story is similar. An A-shift along $x_{1}$ is given by the action of a shift operator $A_{1}$
$A_{1}=2^{-L^{2}/2}\prod_{x_{2}=0}^{L-1}\bigg{(}$ $\prod^{L-1}_{x_{1}=0}$
$\Big{(}1-\xi_{1}(x)\lambda^{1}(x_{1},x_{2})\lambda^{1}(x_{1}+1,x_{2})\Big{)}\bigg{)},$ (61)
with $\xi_{1}(x)=\left(-1\right)^{x_{2}}$. Similarly, a shift along $x_{2}$ is generated by the operator
$A_{2}=2^{-L^{2}/2}\prod_{x_{1}=0}^{L-1}\bigg{(}$ $\prod^{L-1}_{x_{2}=0}$
$\Big{(}1-\xi_{2}(x)\lambda^{1}(x_{1},x_{2})\lambda^{1}(x_{1},x_{2}+1)\Big{)}\bigg{)}.$ (62)
with $\xi_{2}(x)=1$. The B-shifts work in the same way with $\lambda^{1}(x)\to\lambda^{2}(x)$. This allows us to write $\hat{S}_{k}$ in the form
$\hat{S}_{1}$ $=2^{-L^{2}}\prod_{a=1}^{2}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{1}=0}\Big{(}1-\xi_{1}(x)\lambda^{a}(x_{1},x_{2})\lambda^{a}(x_{1}+1,x_{2})\Big{)}\bigg{)}$
$\hat{S}_{2}$ $=2^{-L^{2}}\prod_{a=1}^{2}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{2}=0}\Big{(}1-\xi_{2}(x)\lambda^{a}(x_{1},x_{2})\lambda^{a}(x_{1},x_{2}+1)\Big{)}\bigg{)}$ (63)
When $L$ is a multiple of 4 time reversal can be achieved using
${\cal T}={\cal K}\,\bigg{(}\prod_{x\;{\rm even}}\lambda^{1}(x_{1},x_{2})\bigg{)}\bigg{(}\prod_{x\;{\rm odd}}\lambda^{2}(x_{1},x_{2})\bigg{)},$ (64)
while when $L$ is an odd multiple of 2 we again use $G{\cal T}$ instead. In three dimensions the $\hat{S}$ shifts are
$\hat{S}_{1}$ $=2^{-L^{3}}\prod_{a=1}^{2}\prod_{x_{3}=0}^{L-1}\bigg{(}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{1}=0}$
$\Big{(}1-\xi_{1}(x)\lambda^{a}(x_{1},x_{2},x_{3})\lambda^{a}(x_{1}+1,x_{2},x_{3})\Big{)}\bigg{)}\bigg{)}$
$\hat{S}_{2}$ $=2^{-L^{3}}\prod_{a=1}^{2}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod_{x_{3}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{2}=0}$
$\Big{(}1-\xi_{2}(x)\lambda^{a}(x_{1},x_{2},x_{3})\lambda^{a}(x_{1},x_{2}+1,x_{3})\Big{)}\bigg{)}\bigg{)}$
$\hat{S}_{3}$ $=2^{-L^{3}}\prod_{a=1}^{2}\prod_{x_{2}=0}^{L-1}\bigg{(}\prod_{x_{1}=0}^{L-1}\bigg{(}\prod^{L-1}_{x_{3}=0}$
$\Big{(}1-\xi_{3}(x)\lambda^{a}(x_{1},x_{2},x_{3})\lambda^{a}(x_{1},x_{2},x_{3}+1)\Big{)}\bigg{)}\bigg{)}$ (65)
[…]
${\cal T}={\cal K}\,\bigg{(}\prod_{\sum_{i}x_{i}={\rm even}}\lambda^{1}(x_{1},x_{2},x_{3})\bigg{)}\bigg{(}\prod_{\sum_{i}x_{i}={\rm odd}}\lambda^{2}(x_{1},x_{2},x_{3})\bigg{)}$ (66)
[…]
$R=2^{-L^{d}}\prod_{a=1}^{2}\prod^{L^{d}}_{x=0}\Big{(}1-\frac{1}{2}\epsilon^{ab}\lambda^{a}(x)\lambda^{b}(x)\Big{)}$ (67)
To write down operators that correspond to multiple shifts one simply compounds a series of single shift operators as discussed earlier. As observed earlier, one can verify that all the operators mutually commute with each other except for $S_{k}$ (and indeed all odd shifts) and ${\cal T}$ which anticommute
$\{S_{k},{\cal T}\}=0$ (68)
This fact ensures that any state of the system, including the ground state, has to be at least two fold degenerate (see the appendix for a simple argument) even in the thermodynamic limit. It thus allows for spontaneous symmetry breaking. Furthermore, the definition of $S_{k}$ involves an element $R$ of the $U(1)$ symmetry $M$. Notice that $R$ satisfies $R^{4}=1$ and hence $R$ belongs to a $Z_{4}$ subgroup of $M$. The shifts are clearly symmetries for any element of this $Z_{4}$. The fact that $S_{k}$ and ${\cal T}$ do not commute implies that any attempt to gauge this $Z_{4}$ subgroup will break ${\cal T}$ a mixed lattice ’t Hooft anomaly. In particular, the singlet $U_{A}(1)$ symmetry corresponding to the 3-shift $S_{123}$ will be broken if this subgroup of the vector symmetry is gauged. This is analogous to the result obtained in [9]. In the next section we will explore how these ’t Hooft anomalies can be canceled.
VII Anomaly cancellation
The question we would like to address is whether the mixed ’t Hooft anomaly we found can be canceled. One way to approach this question is to ask whether we can design interactions that can gap the fermions without breaking the $Z_{4}$ symmetry. One way to do the latter is to add $Z_{4}$ invariant four fermion interactions to the theory. The simplest term we can write down takes the form
$G\sum_{x}\chi^{1}(x)\chi^{2}(x)\chi^{3}(x)\chi^{4}(x)+{\rm h.c}$ (69)
which requires four complex staggered fermions and exhibits an explicit $SU(4)=SO(6)$ global symmetry as well as time reversal and shift invariance. For $G\to 0$ one expects the ground state to be eight fold degenerate since it corresponds to eight non-interacting real staggered fermions. However, for $G\to\infty$ the ground state is given by diagonalizing the single site Hamiltonian. It was shown in [25] and [26] that the ground state of this system is in fact a singlet. Indeed, in the latter paper it was shown how to construct a variety of four fermion terms with differing symmetry groups (the maximal symmetry being $SO(7)$) that result in a non-degenerate ground state - generalizing the original result of Kitaev et al [27]. The singlet nature of the ground state implies that the system is incapable of undergoing spontaneous symmetry breaking. This fact in turn implies that the system is free of ’t Hooft anomalies. This phenomenon of producing a gapped, invariant ground state has been termed symmetric mass generation and has already been observed in staggered fermion models with four fermion interactions [27, 19, 16, 28, 29, 30, 18, 31]. Clearly the model can be gapped in this way for eight real staggered fermions each of which yields two Majorana fermions in the naive continuum limit. Thus this analysis confirms the $Z_{16}$ anomaly cancellation condition for Majorana fermions in (3+1) dimensions [32, 33, 26, 34].
VIII Conclusions
In this paper we have examined the shift, time reversal and phase symmetries of Hamiltonian staggered fermions on finite spatial lattices focusing on the case of 3+1 dimensions. We have shown how the shift symmetries correspond to a discrete subgroup of the product of the continuum axial-flavor symmetry and translations. In particular, the odd shifts correspond to axial transformations in the continuum theory and are only symmetries for the massless theory. Furthermore, we find that the odd shifts anticommute with time reversal. We have constructed explicit operators to generate these symmetries along the lines of [14]. The presence of anticommuting symmetries hints at the presence of ’t Hooft anomalies. However, we have argued that the system can be gapped, and hence the ’t Hooft anomaly canceled, for four complex or eight real staggered fields. Indeed, it appears that these mixed shift-time reversal anomalies in the Hamiltonian formalism correspond to the gravitational anomalies of the Euclidean theory.
While writing this paper we became aware of another recent work which also elucidates the symmetry structure of Hamiltonian staggered fermions with the goal of classifying the possibilities for symmetric mass generation [35]. Our results are consistent with their conclusions where the two papers overlap.
Acknowledgements
This work was supported by the US Department of Energy (DOE), Office of Science, Office of High Energy Physics, under Award Number DE-SC0009998. SMC would like to acknowledge useful conversations with Theo Jacobson and Thomas Dumitrescu.
Appendix A Energy level degeneracy
Consider a system which is invariant under time reversal ${\cal T}$ and another symmetry $A$ with $A$ and ${\cal T}$ anticommuting. Thus
$[H,A]=0\quad[H,{\cal T}]=0\quad{\rm with}\quad\{A,{\cal T}\}=0$ (70)
The eigenvectors of $H$ can be labeled with say the eigenvalues $\lambda_{A}$ of $A$
$H\ket{\lambda_{A}}=E\ket{\lambda_{A}}$ (71)
Acting on the left with ${\cal T}$ we find that
$H\,{\cal T}\ket{\lambda_{A}}=E\,{\cal T}\ket{\lambda_{A}}$ (72)
Thus the state ${\cal T}\ket{\lambda_{A}}$ also has energy $E$. To show that it is not proportional to |$\ket{\lambda_{A}}$ consider
$\bra{\lambda_{A}}(A{\cal T}+{\cal T}A)\ket{\lambda_{A}}=0$ (73)
But this implies
$\lambda_{A}\braket{\lambda_{A}}{{\cal T}\lambda_{A}}=0$ (74)
Hence |$\ket{\lambda_{A}}$ and ${\cal T}\ket{\lambda_{A}}$ are orthogonal for $\lambda_{A}\neq 0$.